# Physics-informed Temporal Alignment for Auto-regressive PDE Foundation Models

### v1

@@ Abstract
Auto-regressive partial differential equation (PDE) foundation models have shown great potential in handling time-dependent data. However, these models suffer from the shortcut problem deeply rooted in auto-regressive prediction, causing error accumulation. The challenge becomes particularly evident for out-of-distribution data, as the pretraining performance may approach random model initialization for downstream tasks with long-term dynamics. To deal with this problem, we propose physics-informed temporal alignment (PITA), a self-supervised learning framework inspired by inverse problem solving. Specifically, PITA aligns the physical dynamics discovered at different time steps on each given PDE trajectory by integrating physics-informed constraints into the self-supervision signal. The alignment is derived from observation data without relying on known physics priors, indicating strong generalization ability to the out-of-distribution data. Extensive experiments show that PITA significantly enhances the accuracy and robustness of existing foundation models on diverse time-dependent PDE data. The code is available at https://github.com/SCAILab-USTC/PITA.
@@ 1 Introduction
With the ongoing advancement in computational capabilities and data-driven methodologies (Cicirello, 2024; Brunton & Kutz, 2024), PDE foundation models (Shen et al., 2024; Song et al., 2024b; Gupta & Brandstetter, 2022) have gained increasing prominence in scientific machine learning. These foundational models leverage neural operators as surrogate models to perform PDE trajectory prediction, which runs faster than the traditional numerical solvers. Among the various PDE foundation models, auto-regressive prediction, i.e., predicting the next behavior based on past dynamics data, is one of the most promising strategies for training and pretraining, which aims to endow the models with generalization ability to downstream tasks (Geneva & Zabaras, 2020). However, it is observed that auto-regressive prediction introduces a deeply-rooted shortcut problem, where the model may choose a simple and mendacious solution to approximately satisfy the optimization objective by duplicating previous dynamics. Since the duplicating operation is easy for neural networks to learn, the model may overlook capturing long-term dynamics, resulting in error accumulation. Therefore, the shortcut problem of auto-regressive PDE foundation models remains a significant challenge when facing downstream tasks with long trajectory data.
As we know, each trajectory of PDE data is simulated based on the underlying physical dynamics. Thus, the unique underlying physical equation can govern any segment in the same PDE trajectory. This property naturally facilitates embedding physical constraints into the PDE trajectory prediction. Early researches such as (Raissi et al., 2019; Cuomo et al., 2022) utilize known physics information to construct optimization objectives, enabling precise solutions for specific PDE trajectories. However, this approach may overfit the training data with known physics laws, preventing the models from serving as a foundation for downstream tasks. To this end, some foundation models (Song et al., 2024b; Cao et al., 2024; Zhou et al., 2024b; Sun et al., 2024) introduce PDE symbols as a condition prior to encoding into embedded representations by neural networks, thus guiding the prediction results. Although these methods flexibly encode various known PDE symbols to impose physical constraints on the prediction results, they merely perform conditional mapping based on known equation priors rather than capturing the intrinsic dynamics that represent PDE trajectory data. They may encounter significant performance degradation for downstream tasks without known physical priors. Therefore, exploring universal physical constraints is critically important for PDE foundation models.
@@ 2.1 PDE Foundation Models
2.2 Shortcut Problem
Shortcuts are decision rules that enable models to perform well on standard benchmarks but frequently fail to generalize to more complex testing conditions. This mismatch highlights a gap between the intended and learned solutions (Geirhos et al., 2020). Several frameworks have been proposed to address shortcut learning. For example, the LTGR framework (Du et al., 2021) prevents overconfident predictions on shortcut samples, while COMI (Zhao et al., 2024) reduces the model’s reliance on shortcuts by integrating standard empirical risk minimization to enhance its ability to extract underlying information. The ITSA approach (Chuah et al., 2022) automatically restricts shortcut-related information from being encoded into feature representations. DiffDiv (Scimeca et al., 2023) mitigates bias through diversification using diffusion counterfactuals. Additionally, diffusion-based and auto-regressive generative classifiers have been proposed to address shortcut issues by modeling both causal and spurious features (Li et al., 2024a). Other works (Luo et al., 2021; Brown et al., 2023; Du et al., 2021; Dagaev et al., 2023a; Robinson et al., 2021; Hermann et al., 2023) have also explored various strategies to detect and rectify shortcut learning.
@@ 3.1 Overview
The auto-regressive prediction, detailed in Sec. 3.2, enables iterative future state predictions by progressively leveraging prior outputs. To mitigate shortcut issues inherent in the auto-regressive approach, PDE discovery is incorporated as discussed in Sec. 3.3. This integration encompasses several key steps, including compressing data across the full spatiotemporal domain, constructing tailored function libraries, and employing sparse regression to identify dominant dynamics. Finally, the loss functions comprising data loss, physics loss, and consistency loss are introduced in Sec. 3.4. Together, these losses guide the model to balance among learning from observed data, enforcing physical constraints, and ensuring temporal coherence.
@@ 3.2 Auto-regressive Prediction
where $\boldsymbol{{u}}_{<t}$ represents $\{\boldsymbol{u}_{i}\}_{i=t-T_{in}+1}^{t}$. When $t=T_{in}$, the input sequence consists of the ground truth frames, ensuring that the model starts with accurate information. $\bm{\theta}$ denotes the parameters of the neural operator. To predict the $(t+2)$-th frame, the neural operator uses $T_{in}$ frames along with the predicted frame $\hat{\boldsymbol{u}}_{t+1}$ as input. By iteratively performing this step, the model rolls out a time window of length $T_{ar}$, which can be customized. In our experiments, the roll-in window length is set to $T_{{in}}=10$, and the roll-out window length is set to $T_{ar}=1$ or $T_{ar}=10$. However, directly applying this strategy may lead to the shortcut problem caused by the accumulation of errors propagated across the time window. Our framework employs a physics-informed temporal alignment strategy to address this issue.
@@ 3.3 Governing Equations Discovery
Downsampling Data: Before obtaining the PDE coefficients, the data generated through the auto-regressive prediction is downsampled to reduce computational time. For each $\boldsymbol{u}_{t}$, we assume that the function $\boldsymbol{u}_{t}$ is discretized over $n$ points, represented as ${X}_{n}=\left\{\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{n}\right\}$, $\bm{x}_{i}\in\mathbb{R}^{d}$. In the spatial domain, the grid is randomly sampled to reduce the number of grid points to one-quarter of its original size. The sampled grid is denoted as $\tilde{{X}}_{w}=\{\bm{x}_{i_{1}},\bm{x}_{i_{2}},\cdots,\bm{x}_{i_{w}}\}$, where $w=\lceil n/4\rceil$ and $\left\{i_{1},\cdots,i_{w}\right\}\subseteq\{1,2,\cdots,n\}$ represents the set of sampled indices. The selection of $w$ will be discussed in Sec. 4.4. In the temporal domain, the most recent $T_{C}$ frames are retained, enabling the capture of critical temporal information that is more relevant than the entire sequence. As shown in Figure 2, the diagonal matrix $D$ signifies the downsampling process, with $D_{i,j}\in\{0,1\}$ for $i=j$ and $D_{ij}=0$ for $i\neq j$, where non-zero diagonal elements indicate retaining data. For simplicity, the matrix $D$ is omitted in the following contexts.
@@
3.4 Physics-informed Constraints
The total loss function integrates three parts: the data loss $\mathcal{L}_{Data}$, the physics loss $\mathcal{L}_{Phy}$ and consistency loss $\mathcal{L}_{Con}$.
@@ 3.4 Physics-informed Constraints
Table 1: Comparison of PITA and existing auto-regressive strategies with $T_{ar}=1$. The underlined results represent the predictions of PITA, while the bold results indicate the optimal outcomes for the same model. ’-’ means that the result is unavailable. Notably, five datasets exhibit long trajectory characteristics, with detailed descriptions provided in Appendix B.
@@ 4 Experiments
Datasets: To enable comparisons with baselines employing the auto-regressive strategy, we select 12 datasets from four different sources: 3 datasets from FNO (Li et al., 2020b), 6 datasets from PDEBench (Takamoto et al., 2022), 2 datasets from PDEArena (Gupta & Brandstetter, 2022), and 1 dataset from CFDBench (Luo et al., 2023). For generalization tasks, the Burgers’ equation from (Boussif et al., 2022) is included as an additional dataset. Additional details regarding the datasets used can be found in the Appendix B.
Baselines: We compare PITA with auto-regressive baselines, primarily focusing on PDE foundation models such as DPOT (Hao et al., 2024) and MPP (McCabe et al., 2023). For DPOT, we select models of sizes Ti, S, M, and L, which have been pretrained on 12 datasets. In the case of MPP, we choose the Ti and S models, pretrained on 10 datasets. Both models are designed to work predominantly with 2D datasets. Additionally, we assess our method on single-family models (Shen et al., 2024), where pretraining and finetuning are conducted on a single dataset instead of multiple datasets. Here, FNO-M (Li et al., 2020b) serves as the representative baseline. The sizes and parameters of the various models are detailed in Appendix C.2.
@@ 4.1 State-of-the-Art Results
Overall, PITA reduces the total accumulated error significantly compared to the auto-regressive approach, achieving an improvement of 30.22% over the long trajectory, which translates to an average decrease of 2.00% in error for each time step. The shortcut problem is particularly pronounced in the FNO-NS-1e-4 dataset, where the average percentage of error accumulation per prediction step reaches 34.31%. In contrast, the shortcut problem is less significant in the PDEBench-SWE and PDEBench-DR datasets that are longest trajectory datsets (91 steps). This discrepancy can be attributed to a main factor that the dynamics exhibits smaller variations in long term trajectory, making it easier for the model to capture the temporal dynamics (Takamoto et al., 2022). Despite this, PITA still demonstrates meaningful performance improvements over all the baselines. We further analyze the relevance between the performance gains of PITA and foundation model size. The experiments are conducted on PDEBench-DR with the longest trajectory, as shown in Figure 3. We can see that in long trajectory predictions, PITA demonstrates scalability. It even outperforms the baseline with 100M parameters when using only 30M parameters, showcasing parameter efficiency in modeling long-term dynamics. Additional scaling experiments on long trajectory datasets are presented in Appendix D.3.
@@ 4.3 Generalizing to Downstream tasks
Table 2: Generalization on two downstream tasks of PITA compared with baselines. More comprehensive generalization experiments are presented in Appendix D.4.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (NSFC) under Grant 12425113; the Natural Science Foundation of Jiangsu Province under Grant BK20240462; the China Postdoctoral Science Foundation under Grants 2023M733427 and 2023TQ0349; and the Jiangsu Funding Program for Excellent Postdoctoral Talent. We also gratefully acknowledge the support of the Key Laboratory of the Ministry of Education for Mathematical Foundations and Applications of Digital Technology, USTC, and the valuable discussions and assistance provided by Zijun Ding.
@@ Appendix A Table of Notations
Appendix B Details of Datasets
B.1 FNO-NS-$\nu$
@@ B.2 PDEBench-CNS-$(M,\eta)$
$\partial_{t}\rho+\nabla\cdot(\rho\mathbf{v})$ $=0$ (14)
$,\rho\left(\partial_{t}\mathbf{v}+\mathbf{v}\cdot\nabla\mathbf{v}\right)$ $=-\nabla p+\eta\triangle\mathbf{v}+(\zeta+\eta/3)\nabla(\nabla\cdot\mathbf{v})$

### v2

@@ Abstract
Auto-regressive partial differential equation (PDE) foundation models have shown great potential in handling time-dependent data. However, these models suffer from error accumulation caused by the shortcut problem deeply rooted in auto-regressive prediction. The challenge becomes particularly evident for out-of-distribution data, as the pretraining performance may approach random model initialization for downstream tasks with long-term dynamics. To deal with this problem, we propose physics-informed temporal alignment (PITA), a self-supervised learning framework inspired by inverse problem solving. Specifically, PITA aligns the physical dynamics discovered at different time steps on each given PDE trajectory by integrating physics-informed constraints into the self-supervision signal. The alignment is derived from observation data without relying on known physics priors, indicating strong generalization ability to out-of-distribution data. Extensive experiments show that PITA significantly enhances the accuracy and robustness of existing foundation models on diverse time-dependent PDE data. The code is available at https://github.com/SCAILab-USTC/PITA.
@@ 1 Introduction
With the ongoing advancement in computational capabilities and data-driven methodologies (Cicirello, 2024; Brunton & Kutz, 2024), PDE foundation models (Shen et al., 2024; Song et al., 2024b; Gupta & Brandstetter, 2022) have gained increasing prominence in scientific machine learning. These foundational models leverage neural operators as surrogate models to perform PDE trajectory prediction, which runs significantly faster than the traditional numerical solvers. Among the various PDE foundation models, auto-regressive prediction, i.e., predicting the next behavior based on past dynamics data, is one of the most promising strategies for training or pretraining, which aims to endow the models with generalization ability to downstream tasks (Geneva & Zabaras, 2020). However, it is observed that auto-regressive prediction may introduce a deeply-rooted shortcut problem, where the model chooses a simple and mendacious solution to approximately satisfy the optimization objective by duplicating previous dynamics. More details are discussed in Appendix B. Since the duplicating operation is easy for neural networks to learn, the model may overlook capturing long-term dynamics, resulting in error accumulation. Therefore, the long-term accurate prediction of auto-regressive PDE foundation models remains a significant challenge when facing downstream tasks with long trajectory data.
As we know, all trajectories of PDE data are simulated based on the underlying physical dynamics. Thus, the unique underlying physical equation can govern any segment in the corresponding PDE trajectory. This property naturally facilitates embedding physical constraints into the PDE trajectory prediction. Early research such as (Raissi et al., 2019; Cuomo et al., 2022) utilizes known physics information to construct optimization objectives, enabling precise solutions for specific PDE trajectories. However, this approach may overfit the training data with known physics laws, preventing the models from serving as a foundation for downstream tasks. To this end, some foundation models (Song et al., 2024b; Cao et al., 2024; Zhou et al., 2024b; Sun et al., 2024) introduce PDE symbols as a condition prior to encoding into embedded representations by neural networks, thus guiding the prediction results. Although these methods flexibly encode various known PDE symbols to impose physical constraints on the prediction results, they merely perform conditional mapping based on known equation priors rather than capturing the intrinsic dynamics that represent PDE trajectory data. They may encounter significant performance degradation for downstream tasks without known physical priors. Therefore, exploring universal physical constraints is critically important for PDE foundation models.
@@ 2.1 PDE Foundation Models
2.2 Shortcut Problem and Error Accumulation
Shortcuts commonly emerge as decision rules allowing models to excel on standard benchmarks while struggling to generalize under more complex testing conditions. This discrepancy between intended and learned solutions (Geirhos et al., 2020) leads to severe error accumulation, particularly in long temporal sequences. Several frameworks have been proposed to address shortcut learning. For example, the LTGR framework (Du et al., 2021) prevents overconfident predictions on shortcut samples, while COMI (Zhao et al., 2024) reduces the model’s reliance on shortcuts by integrating standard empirical risk minimization to enhance its ability to extract underlying information. Additionally, diffusion-based and auto-regressive generative classifiers have been proposed to address shortcut issues by modeling both causal and spurious features (Li et al., 2024a). Other works (Luo et al., 2021; Brown et al., 2023; Du et al., 2021; Dagaev et al., 2023a; Robinson et al., 2021; Hermann et al., 2023; Scimeca et al., 2023; Chuah et al., 2022) have also explored various strategies to detect and rectify shortcut learning.
@@ 3.1 Overview
The auto-regressive prediction, detailed in Sec. 3.2, enables iterative future state predictions by progressively leveraging prior outputs. To mitigate error accumulation inherent in the auto-regressive approach, PDE discovery is incorporated as discussed in Sec. 3.3. This integration encompasses several key steps, including compressing data across the full spatiotemporal domain, constructing tailored function libraries, and employing sparse regression to identify dominant dynamics. Finally, the loss functions comprising data loss, physics loss, and consistency loss are introduced in Sec. 3.4. Together, these losses guide the model to balance among learning from observed data, enforcing physical constraints, and ensuring temporal coherence.
@@ 3.2 Auto-regressive Prediction
where $\boldsymbol{{u}}_{<t}$ represents $\{\boldsymbol{u}_{i}\}_{i=t-T_{in}+1}^{t}$. When $t=T_{in}$, the input sequence denotes the ground truth frames, ensuring that the model starts with accurate information. $\bm{\theta}$ denotes the parameters of the neural operator. To predict the $(t+2)$-th frame, the neural operator uses $T_{in}$ frames along with the predicted frame $\hat{\boldsymbol{u}}_{t+1}$ as input. By iteratively performing this step, the model rolls out a time window of length $T_{ar}$, which can be customized. In our experiments, the roll-in window length is set to $T_{{in}}=10$, and the roll-out window length is set to $T_{ar}=1$ or $T_{ar}=10$. However, directly applying this strategy may lead to the shortcut problem, causing the accumulation of errors propagated across the time window. Our framework employs a physics-informed temporal alignment strategy to address this issue.
@@ 3.3 Governing Equations Discovery
Downsampling Data: Before obtaining the PDE coefficients, the data generated through the auto-regressive prediction is downsampled to reduce computational time. For each $\boldsymbol{u}_{t}$, we assume that the function $\boldsymbol{u}_{t}$ is discretized over $n$ points, represented as ${X}_{n}=\left\{\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{n}\right\}$, $\bm{x}_{i}\in\mathbb{R}^{d}$. In the spatial domain, the grid is randomly sampled to reduce the number of points to one-quarter of its original size. The sampled grid is denoted as $\tilde{{X}}_{w}=\{\bm{x}_{i_{1}},\bm{x}_{i_{2}},\cdots,\bm{x}_{i_{w}}\}$, where $w=\lceil n/4\rceil$ and $\left\{i_{1},\cdots,i_{w}\right\}\subseteq\{1,2,\cdots,n\}$ represents the set of sampled indices. The selection of $w$ will be discussed in Sec. 4.4. The derivatives are computed using a small number of spatially localized points near each measurement position via polynomial interpolation (Rudy et al., 2017). In the temporal domain, the most recent $T_{C}$ frames are retained, enabling the capture of critical temporal information that is more relevant than the entire sequence. As shown in Figure 2, the diagonal matrix $D$ signifies the downsampling process, with $D_{i,j}\in\{0,1\}$ for $i=j$ and $D_{ij}=0$ for $i\neq j$, where non-zero diagonal elements indicate retaining data. For simplicity, the matrix $D$ is omitted in the following contexts.
@@
3.4 Integrating Data and Physics Supervisions
The total loss function integrates three parts: the data loss $\mathcal{L}_{Data}$, the physics loss $\mathcal{L}_{Phy}$ and the consistency loss $\mathcal{L}_{Con}$.
@@ 3.4 Integrating Data and Physics Supervisions
Table 1: Comparison of PITA and existing auto-regressive strategies with $T_{ar}=1$. The underlined results represent the predictions of PITA, while the bold results indicate the optimal outcomes for the same model. ’-’ means that the result is unavailable. Notably, five datasets exhibit long trajectory characteristics, with detailed descriptions provided in Appendix C.
@@ 4 Experiments
Datasets: To enable comparisons with baselines employing the auto-regressive strategy, we select 12 datasets from four different sources: 3 datasets from FNO (Li et al., 2020b), 6 datasets from PDEBench (Takamoto et al., 2022), 2 datasets from PDEArena (Gupta & Brandstetter, 2022), and 1 dataset from CFDBench (Luo et al., 2023). For generalization tasks, the Burgers’ equation from (Boussif et al., 2022) is included as an additional dataset. Additional details regarding the datasets used can be found in the Appendix C.
Baselines: We compare PITA with auto-regressive baselines, primarily focusing on PDE foundation models such as DPOT (Hao et al., 2024) and MPP (McCabe et al., 2023). For DPOT, we select models of sizes Ti, S, M, and L, which have been pretrained on 12 datasets. In the case of MPP, we choose the Ti and S models, pretrained on 10 datasets. Both models are designed to work predominantly with 2D datasets. Additionally, we assess our method on single-family models (Shen et al., 2024), where pretraining and finetuning are conducted on a single dataset instead of multiple datasets. FNO-M (Li et al., 2020b) serves as the representative baseline. The sizes and parameters of the various models are detailed in Appendix D.1. These ensure a comprehensive and fair evaluation across diverse PDE settings.
@@ 4.1 State-of-the-Art Results
Overall, PITA reduces the total accumulated error significantly compared to the auto-regressive approach, achieving an improvement of 30.22% over the long trajectory, which translates to an average decrease of 2.00% in error for each time step. The shortcut problem is particularly pronounced in the FNO-NS-1e-4 dataset, where the average percentage of error accumulation per prediction step reaches 34.31%. In contrast, the shortcut problem is less significant in the PDEBench-SWE and PDEBench-DR datasets that are longest trajectory datsets (91 steps). This discrepancy can be attributed to a main factor that the dynamics exhibits smaller variations in long term trajectory, making it easier for the model to capture the temporal dynamics (Takamoto et al., 2022). Despite this, PITA still demonstrates meaningful performance improvements over all the baselines. We further analyze the relevance between the performance gains of PITA and foundation model size. The experiments are conducted on PDEBench-DR with the longest trajectory, as shown in Figure 3. We can see that in long trajectory predictions, PITA demonstrates scalability. It even outperforms the baseline with 100M parameters when using only 30M parameters, showcasing parameter efficiency in modeling long-term dynamics. Additional scaling experiments on long trajectory datasets are presented in Appendix E.3.
@@ 4.3 Generalizing to Downstream tasks
Table 2: Generalization on two downstream tasks of PITA compared with baselines. More comprehensive generalization experiments are presented in Appendix E.4.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (NSFC) Major Research Plan on Interpretable and General Purpose Next-generation Artificial Intelligence under Grant No. 92370205, NSFC Grant 12425113, the Natural Science Foundation of Jiangsu Province under Grant BK20240462, the China Postdoctoral Science Foundation under Grants 2023M733427 and 2023TQ0349, and the Jiangsu Funding Program for Excellent Postdoctoral Talent. We also acknowledge support from the Key Laboratory of the Ministry of Education for Mathematical Foundations and Applications of Digital Technology, USTC, and the valuable discussions and assistance provided by Zijun Ding.
@@ Appendix A Table of Notations
Appendix B Illustration of the Shortcut Problem and Error Accumulation
This section investigates the origin of the shortcut problem and its consequential error accumulation in auto-regressive prediction frameworks. When predicting the $T$-th frame, the prediction error at the ($T$-$1$)-th time step propagates to subsequent predictions. This propagation triggers cumulative error amplification over successive time steps, which is a critical limitation of auto-regressive mechanisms. As illustrated in Figure 5, within a prediction window of length $T_{ar}$, errors originating from the initial frame amplify iteratively, thereby degrading the accuracy of later predictions. Such accumulation manifests as severe deviations or unphysical artefacts in dynamical system simulations, particularly when predicted states violate the fundamental physical laws represented by the governing PDEs. For instance, in fluid dynamics or climate modeling applications, these artefacts may manifest as non-physical vorticity patterns or temperature distributions that diverge from observed system behavior. The compounding errors not only distort short-term predictions but also erode the validity of extrapolated trajectories over extended timescales.
Figure 5: Illustration of Error Propagation in Auto-regressive Prediction.
Furthermore, the error accumulation phenomenon may arise from shortcut learning in auto-regressive approaches, where models prioritize copying values from the ($T$-$1$)-th frame to predict the $T$-th frame, rather than learning the underlying physical dynamics governed by PDEs. To quantify this behavior, we analyze two long-trajectory datasets using auto-regressive prediction, where the nRMSE is notably higher compared to other methods, as shown in Figure 6. Where the pink curve labeled Temporal SSIM (predicted $t$ vs predicted $t-1$) measures the structural similarity index (SSIM) between consecutive predicted frames, $\hat{\boldsymbol{u}}_{t-1}$ and $\hat{\boldsymbol{u}}_{t}$. The blue curve labeled Frame SSIM (predicted $t$ vs ground truth $t$ ) quantifies the SSIM between the predicted frame $\hat{\boldsymbol{u}}_{t}$ and the ground truth frame $\boldsymbol{u}_{t}$.
Figure 6: Intrinsic connection between error accumulation and shortcut learning. On the left, we present the shortcut issue on the PDEBench-DR dataset with a trajectory length being 91. On the right, we present the shortcut issue on the NS2D-Force dataset with a trajectory length being 46.
At initial time steps, the blue curve dominates the pink curve (Figure 6), indicating that $\hat{\boldsymbol{u}}_{t}$ aligns more closely with the ground truth $\boldsymbol{u}_{t}$ than with the prior prediction $\hat{\boldsymbol{u}}_{t-1}$. This suggests the model initially captures physical dynamics encoded in the PDEs rather than relying on superficial shortcut features. However, at later stages, the pink curve surpasses the blue curve, revealing that $\hat{\boldsymbol{u}}_{t}$ increasingly resembles $\hat{\boldsymbol{u}}_{t-1}$ instead of $\boldsymbol{u}_{t}$, thereby exacerbating shortcut-driven error propagation. This divergence reflects the model’s failure to learn true physical features instead, it recursively replicates prior predictions, introducing incremental errors at each step. Over extended trajectories, the repeated copying of erroneous values leads to severe error accumulation, particularly in systems requiring long-term stability, such as turbulence modeling or climate projections. Addressing this limitation is critical for improving the robustness of auto-regressive methods in scientific computing.
Appendix C Details of Datasets
C.1 FNO-NS-$\nu$
@@ C.2 PDEBench-CNS-$(M,\eta)$
$\partial_{t}\rho+\nabla\cdot(\rho\mathbf{v})$ $=0,$ (14)
$\rho\left(\partial_{t}\mathbf{v}+\mathbf{v}\cdot\nabla\mathbf{v}\right)$ $=-\nabla p+\eta\triangle\mathbf{v}+(\zeta+\eta/3)\nabla(\nabla\cdot\mathbf{v})$